Rachel has 25% acid solution and a 50% acid solution. How many liters of each solution should she mix if she wants to obtain 100 liters of a 40% acid solution?
Answer by Alan3354(69443)

Rachel has 25% acid solution and a 50% acid solution. How many liters of each solution should she mix if she wants to obtain 100 liters of a 40% acid solution?
----------
t = 25%
f = 50%
---
t + f = 100 ---- total solution
25t + 50f = 100*40 ------- total acid
==========
Solve them.